# Impact of phylogeny on the inference of functional sectors from protein sequence data

### v1

@@ Acknowledgments
A.-F. B. thanks Ned S. Wingreen and Shou-Wen Wang for inspiring discussions. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 851173, to A.-F. B.).
@@
Supplementary material
S1.1 Complex saddle point approximation
$H=\frac{\kappa}{2}\Big{(}\sum_{l}\sigma_{l}D_{l}-\tau^{*}\Big{)}^{2}+\sum_{l}\sigma_{l}f_{l},$ (14)
$Z=\sum_{\{\sigma_{l}\}}\exp(-H)=\sum_{\{\sigma_{l}\}}\exp\Big{(}-\frac{\kappa}{2}\Big{[}\sum_{l}\sigma_{l}D_{l}-\tau^{*}\Big{]}^{2}-\sum_{l}\sigma_{l}f_{l}\Big{)}.$ (15)
$m_{l}=-\left.\frac{\partial\ln Z}{\partial f_{l}}\right|_{\vec{f}=\vec{0}},$ (16)
$C_{ll^{\prime}}=\left.\frac{\partial^{2}\ln Z}{\partial f_{l^{\prime}}\partial f_{l}}\right|_{\vec{f}=\vec{0}}=-\left.\frac{\partial m_{l}}{\partial f_{l^{\prime}}}\right|_{\vec{f}=\vec{0}}.$ (17)
To compute the partition function $Z$, we employ the Hubbard-Stratonovich transform, which involves an auxiliary field $h$:
[…]
Employing this result, we have
[…]
where we introduced
$a(h)\equiv\exp\left[ih\tau^{*}+\sum_{l}\ln\cosh\Big{(}ihD_{l}+f_{l}\Big{)}\right].$ (20)
[…]
$\ln Z=\ln\left[2^{L}a(0)\right]+\frac{\kappa}{2}\frac{a^{(2)}(0)}{a(0)}+\frac{\kappa^{2}}{8}\left\{\frac{a^{(4)}(0)}{a(0)}-\left[\frac{a^{(2)}(0)}{a(0)}\right]^{2}\right\}+O(\kappa^{3})\,.$ (22)
Computing the successive derivatives of $a$ (see Eq 20) and employing the definitions in Eq 16 and 17, we obtain
$m_{k}$ $=\tau^{*}\kappa D_{k}\left[1+\kappa\left(D_{k}^{2}-\sum_{l}D_{l}^{2}\right)\right]+O(\kappa^{3}),$ (23)
$C_{jk}$ $=-\kappa D_{j}D_{k}\left[1+\kappa\left(D_{k}^{2}+D_{j}^{2}-\sum_{l}D_{l}^{2}\right)\right]\left(1-\delta_{jk}\right)+\left(1-{\tau^{*}}^{2}\kappa^{2}D_{k}^{2}\right)\delta_{jk}+O(\kappa^{3}).$ (24)
We also obtain the elements of the inverse correlation matrix:
$C^{-1}_{jk}=\kappa D_{j}D_{k}\left(1-\delta_{jk}\right)+\left[1+\kappa^{2}D_{j}^{2}\left(\sum_{l}D_{l}^{2}-D_{j}^{2}+{\tau^{*}}^{2}\right)\right]\delta_{jk}+O(\kappa^{3}).$ (26)
$\tilde{C}^{-1}_{jk}=\kappa D_{j}D_{k}\left(1-\delta_{jk}\right)+O(\kappa^{3})\,,$ (27)
S1.2 High-temperature Plefka expansion of the TAP free energy
We can derive the expression of the inverse covariance matrix in the large temperature limit (or equivalently here, in the limit of small coupling $\kappa$), using a Plefka expansion. We follow the calculation from [75], starting from the Hamiltonian in Eq 14 with no auxiliary field, and ignoring additive constant terms:
$H$ $=\dfrac{\kappa}{2}\sum_{i,j}D_{i}D_{j}\sigma_{i}\sigma_{j}-\kappa\sum_{j}\tau^{*}\sigma_{j}D_{j}$ (28)
$=\dfrac{\tilde{\kappa}}{2}\sum_{i,j}\tilde{D}_{i}\tilde{D}_{j}\sigma_{i}\sigma_{j}-\tilde{\kappa}\sum_{j}\tilde{\tau}\sigma_{j}\tilde{D_{j}},$ (29)
where we defined rescaled parameters as
$\tilde{\kappa}$ $=\kappa\sum_{j}D_{j}^{2},$ (30)
$\tilde{\tau}$ $=\dfrac{\tau^{*}}{\sqrt{\sum_{j}D_{j}^{2}}},$ (31)
$\tilde{D}_{i}$ $=\dfrac{D_{i}}{\sqrt{\sum_{j}D_{j}^{2}}}.$ (32)
S1.2.1 TAP free energy expansion
$\Phi(\tilde{\kappa})=\Phi(0)+\dfrac{\tilde{\kappa}}{2}\sum_{i,j}J_{ij}m_{i}m_{j}-\tilde{\kappa}\sum_{i}h_{i}m_{i}+\sum_{p=1}^{4}\dfrac{\tilde{\kappa}^{p}}{2p}\sum_{\begin{subarray}{c}i_{1},...,i_{p}\\ \text{pairwise distinct}\end{subarray}}J_{i_{1}i_{2}}...J_{i_{p}i_{1}}\prod_{\alpha=1}^{p}v_{i_{\alpha}},$ (33)
where
$\Phi(0)=-\sum_{i}\left[\dfrac{1-m_{i}}{2}\log\left(\dfrac{1-m_{i}}{2}\right)+\dfrac{1+m_{i}}{2}\log\left(\dfrac{1+m_{i}}{2}\right)\right]\,.$ (34)
S1.2.2 Extremized magnetizations
$\frac{\partial\Phi}{\partial m_{i}}=\dfrac{1}{2}\log\dfrac{1+m_{i}}{1-m_{i}}+\tilde{\kappa}\sum_{j}J_{ij}m_{j}-\tilde{\kappa}h_{i}-\tilde{\kappa}^{2}\sum_{\begin{subarray}{c}j\\ j\neq i\end{subarray}}J_{ij}^{2}(1-m_{j}^{2})m_{i}\\ -\tilde{\kappa}^{3}\sum_{\begin{subarray}{c}j\neq k\\ j,k\neq i\end{subarray}}J_{ij}J_{jk}J_{ki}(1-m_{j}^{2})(1-m_{k}^{2})m_{i}-\tilde{\kappa}^{4}\sum_{\begin{subarray}{c}j\neq k\neq l\\ j,k,l\neq i\end{subarray}}J_{ij}J_{jk}J_{kl}J_{li}(1-m_{j}^{2})(1-m_{k}^{2})(1-m_{l}^{2})m_{i}=0.$ (35)
We write the perturbative expansion $m_{i}=m_{i}^{(0)}+\tilde{\kappa}m_{i}^{(1)}+\tilde{\kappa}^{2}m_{i}^{(2)}+\tilde{\kappa}^{3}m_{i}^{(3)}+\tilde{\kappa}^{4}m_{i}^{(4)}$, where we only keep leading terms in $N$ (in the large $N$ limit) in all $m_{i}^{(k)}$. Inserting this expression in 35, we obtain
$m_{i}^{(0)}$ $=0$ (36a)
$m_{i}^{(1)}$ $=-h_{i}=\tilde{\tau}\tilde{D}_{i}$ (36b)
$m_{i}^{(2)}$ $=\sum_{j}J_{ij}m_{j}^{(1)}=-\tilde{\tau}\tilde{D}_{i}\sum_{\begin{subarray}{c}j\end{subarray}}\tilde{D}_{j}^{2}$ (36c)
$m_{i}^{(3)}$ $=\sum_{j}J_{ij}m_{j}^{(2)}=\tilde{\tau}\tilde{D}_{i}\left(\sum_{\begin{subarray}{c}j\end{subarray}}\tilde{D}_{j}^{2}\right)^{2}$ (36d)
$m_{i}^{(4)}$ $=\sum_{j}J_{ij}m_{j}^{(3)}=-\tilde{\tau}\tilde{D}_{i}\left(\sum_{\begin{subarray}{c}j\end{subarray}}\tilde{D}_{j}^{2}\right)^{3}.$ (36e)
Note that the dominant terms we retained are of order $1/\sqrt{N}$.
S1.2.3 Inverse covariance matrix
From the free energy, we can derive the elements of the inverse covariance matrix through $C^{-1}_{ij}=-\dfrac{\partial^{2}\Phi}{\partial m_{i}\partial m_{j}}$. Differentiating Eq 35 yields
$+\tilde{\kappa}^{4}\sum_{\begin{subarray}{c}j\neq k\neq l\\ j,k,l\neq i\end{subarray}}J_{ij}J_{jk}J_{kl}J_{li}(1-m_{j}^{2})(1-m_{k}^{2})(1-m_{l}^{2})+O(\tilde{\kappa}^{5}),$ (37)
$-6\tilde{\kappa}^{4}\sum_{\begin{subarray}{c}k\neq l\\ k,l\neq i,j\end{subarray}}J_{ij}J_{jk}J_{kl}J_{li}(1-m_{k}^{2})(1-m_{l}^{2})m_{i}m_{j}+O(\tilde{\kappa}^{5})\,\,\,\,\textrm{if $i\neq j$}.$ (38)
Injecting Eq 36 finally yields, up to order 4 in $\tilde{\kappa}$:
$+\tilde{\kappa}^{4}\tilde{D}_{i}^{2}\left(\tilde{\tau}^{4}\tilde{D}_{i}^{2}+3\tilde{\tau}^{2}(\sum_{\begin{subarray}{c}k\end{subarray}}\tilde{D}_{k}^{2})^{2}+\sum_{\begin{subarray}{c}j\neq k\neq l\\ j,k,l\neq i\end{subarray}}\tilde{D}_{j}^{2}\tilde{D}_{k}^{2}\tilde{D}_{l}^{2}\right)+O(\tilde{\kappa}^{5})\,,$ (39)
$C_{ij}^{-1}=$ $\tilde{\kappa}\tilde{D}_{i}\tilde{D}_{j}+O(\tilde{\kappa}^{5})\,\,\,\,\textrm{if $i\neq j$}.$ (40)
Except the $1$ in the first line, all other terms are of order $1/N$. These expressions are consistent with the second order expansion in $\kappa$ of the inverse correlation matrix obtained in Eq 26. They further extend our previous results on the ICOD matrix [31], up to order 4 in $\tilde{\kappa}$.
S1.3 Comparison to synthetic data
In Fig S1, we compare our analytical approximations from Eq S1.2.3, 40 to the elements of the inverse covariance matrix obtained directly from synthetic data. We find a good agreement, which improves when the number $M$ of synthetic sequences increases, and, in the case of the diagonal elements, with the order of the expansion in $\tilde{\kappa}$. Despite the fact that these expansions are made in the limit of small selection strength $\tilde{\kappa}$, when $\tilde{\kappa}$ is very small, the data is noisy, and agreement is thus less good that when $\tilde{\kappa}$ is somewhat larger. Interestingly, the analytical approximation Eq 40 for off-diagonal elements is very good even for intermediate and large values of $\tilde{\kappa}$, hinting that its validity extends beyond the regime of small $\tilde{\kappa}$ and beyond order 4 in $\tilde{\kappa}$. Meanwhile, for the diagonal terms, the expansion gives good results for relatively small $\tilde{\kappa}$, but agreement becomes less good as $\tilde{\kappa}$ becomes larger, in agreement with expectations for a small-$\tilde{\kappa}$ expansion.
Figure S1: Inverse covariance matrix: Comparison between analytical approximation and synthetic data. Left (rep. right) panels show the slope of the linear fit and the Pearson correlation between off-diagonal (resp. diagonal) elements of the inverse covariance matrix given by analytical formulas (Eq S1.2.3, 40) and the inverse covariance matrix computed from synthetic data. For the latter, we generate independent sequences at equilibrium as in Fig 2 but the parameter $\kappa$ is varied and $\tau^{*}$ is taken of order 1 (specifically, $\tau^{*}=0.05\sqrt{\sum_{i}D_{i}^{2}}\approx 1$). In the left panels, we construct MSAs with depths $M=10000,M=30000,M=50000$. In the right panels, we take $M=50000$ sequences and compare the diagonal of the inverse covariance matrix computed from the data to the analytical formula S1.2.3 by gradually adding orders in $\tilde{\kappa}$. Note that we do not use a pseudocount ($a=0$) here, which is acceptable given the large numbers of sequences, and allows a more direct comparison with the analytical formulas. The dashed lines represent the ideal values (slope 1 and Pearson correlation 1), which would indicate a perfect match between the data and the analytical approximation.
S2 Supplementary figures
@@
Figure S3: Impact of phylogeny on mutational effect recovery for SCA. Same as in Fig 4, but for SCA. For SCA, eigenvectors associated to the largest ($\lambda_{max}$) and smallest ($\lambda_{min}$) eigenvalues are considered. The results for ICOD and conservation (see Fig 4) are reproduced here for comparison purposes.
Figure S5: Impact of the favored trait value on mutational effect recovery at the opposite end of the spectrum. Same as in Fig 5, but focusing on eigenvectors at the opposite end of the spectrum. For ICOD (resp. SCA), eigenvectors associated to the smallest eigenvalue $\Lambda_{min}$ and (resp. $\lambda_{min}$) are used, while for covariance, the eigenvector associated to the largest eigenvalue $\lambda_{max}$ is used. Conservation results (see Fig 5) are reproduced here for comparison purposes.
Figure S6: Impact of the favored trait value on symmetrized AUC. Same as in Fig 5, but instead of using recovery, we consider the symmetrized AUC between the eigenvectors and the mutational effects.
Figure S7: Impact of the favored trait value on symmetrized AUC at the opposite end of the spectrum. Same as in Fig S5, but instead of using recovery, we consider the symmetrized AUC between the eigenvectors and the mutational effects.
Figure S8: Impact of selection strength on mutational effect recovery for SCA and conservation. Same as in Fig 6 but using SCA and conservation. For SCA, the eigenvector associated to the largest eigenvalue ($\lambda_{max}$) is used. ICOD results (see Fig 6) are reproduced here for comparison purposes.
Figure S9: Impact of large favored trait values on ICOD spectrum and eigenvectors. Top panels show the eigenvalues of the ICOD matrix at different values of favored trait value $\tau^{*}$ (left: $\tau^{*}=90$; right: $\tau^{*}=195$ and $\tau^{*}=300$). Inset in the upper right panel: zoom on the 20 first eigenvalues. Bottom panels show the normalised eigenvectors associated to the largest eigenvalue ($\Lambda_{max}$) on the left and smallest eigenvalue ($\Lambda_{min}$) on the right panel, for the same three values of $\tau^{*}$ as in the top panels. In the bottom panels, we only show the first 50 sites of the eigenvectors, which comprise the 20 sites associated to the sector and 30 non-sector sites. The vector of mutational effects $\vec{D}$ is normalised and shown for comparison with the eigenvectors. Data is generated as in the equilibrium (no phylogeny) case in Fig 3, but using different values of $\tau^{*}$.

### v2

@@ Acknowledgments
A.-F. B. thanks Ned S. Wingreen and Shou-Wen Wang for inspiring discussions. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 851173, to A.-F. B.). The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
@@
Supplementary files
Supplementary Appendix.
The Supplementary Appendix, included at the end of the present document, comprises calculations for the analytical approximation of the inverse covariance matrix at higher orders (section S1) and an analysis of the impact of selection parameters on mutational effect recovery (section S2).
Figure S2: Spectrum of the ICOD matrix and of its block diagonal approximation. Same as in Fig 2, except that the vector $\vec{D}$ of mutational effects comprises both negative and positive components. Specifically, we took the same $\vec{D}$ as in Fig 2, but we multiplied the 10 first components (corresponding to half of the sector) by $-1$.
Figure S3: Impact of phylogeny on mutational effect recovery by the opposite end of the spectrum. Same as in Fig 4, but using the eigenvectors associated to the eigenvalues at the opposite end of the spectrum. For ICOD (resp. SCA), eigenvectors associated to the smallest eigenvalue $\Lambda_{min}$ (resp. $\lambda_{min}$) are considered. For covariance $C$, the eigenvector associated to the largest eigenvalue $\lambda_{max}$ is considered.
@@
Figure S5: Phylogenetic trees in the Beta-coalescent process. Beta-coalescent trees are generated using the code associated to Ref. [49], which is available on GitHub (https://github.com/rneher/betatree). Each column corresponds to a specific value of the parameter $\alpha$ which characterizes the Beta-coalescent tree. For each value of $\alpha$, we sampled two trees, represented in the two rows of the figure. Note that $\alpha=1$ corresponds to the Bolthausen-Sznitman coalescent, while $\alpha=2$ corresponds to the Kingman coalescent.
Figure S6: Impact of phylogeny on mutational effect recovery for different phylogenetic trees. Same analysis as in Fig 4, but using data generated along the Beta-coalescent trees shown in Fig S5 instead of a perfect binary tree. Each panel here is associated to the tree shown in the corresponding panel of Fig S5. Mutational effect recovery is shown for various methods versus the average number $\langle m\rangle$ of accepted mutations between the ancestor and a leaf. For synthetic data generation, the number of accepted mutations along each branch is drawn from a Poisson law with mean equal to a multiplicative constant times the branch length shown in Fig S5. The multiplicative constant (which is the same for all branches in a tree) is then varied to tune the amount of phylogeny in the data, yielding various values of $\langle m\rangle$. All curves are averaged over 100 realisations of data generation along the same tree. We checked that the asymptotic values obtained without phylogeny (dashed lines) are reached for large $\langle m\rangle$.
@@
Figure S8: Identifying functionally important sites in natural protein families: Mean performance over phylogenetic cutoffs. Same as in Fig 6, but using the mean of the performance, i.e. mean of the symmetrized AUCs, over all phylogenetic cutoffs considered. The error bars represent the standard deviation over phylogenetic cutoffs. Here, the average of the symmetrized AUC over all families is 0.42 for conservation, 0.39 for ICOD, 0.25 for MI and 0.19 for SCA. The mapping between protein family number and name is given in Table S1.
Figure S10: Relative difference of ICOD and SCA performances versus MSA diversity. The relative difference between the symmetrized AUC score $S$ for ICOD and SCA (i.e., $\left(S_{\textrm{ICOD}}-S_{\textrm{SCA}}\right)/S_{\textrm{ICOD}}$) is plotted versus the mean pairwise Hamming distance in the MSA, for the prediction of sites with large mutational effects in 30 different protein families. The red dashed line separates the cases where ICOD is best (positive values) from those where SCA is best (negative values). Performances measured as symmetrized AUC values are the same as in Fig 6, and protein families are listed in Table S1. For each family, the mean pairwise Hamming distance is computed for the MSA associated to the maximum phylogenetic cutoff $C_{max}$ considered, thus reflecting the largest diversity in the family, see Table S2. Phylogenetic cutoffs are defined and their values are given in the paragraph “MSA construction” of the Methods section.